Circular No. 8898

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V598 PUPPIS N. N. Samus, Institute of Astronomy, Russian Academy of Sciences,

informs the Central Bureau that the GCVS designation V598 Pup has been assigned to the apparent nova initially posted by Read et al. (http://www.

astronomerstelegram.org/?read=1282) as a bright x-ray transient (as

seen with the XMM-Newton satellite on Oct. 8 at 0.2–2 keV) that is located at α = 7^{h}05^{m}42^{s}.7, δ = −38^{o}14′42′′ (equinox 2000.0), and subsequently found as a bright star optically (within the reported 8′′ x-ray error circle) by Torres et al. (ibid., /?read=1285), who explain that their spectroscopy obtained on Nov. 16.34 UT reveal numerous features suggestive of a nova in the auroral phase. E. O. Waagen, AAVSO, writes that S. Dvorak (Clermont, FL, U.S.A.) reports V = 10.28 and B = 11.25 for V598 Pup from CCD images obtained on Nov. 18.38; A. Henden, AAVSO, has measured the following position end figures from Dvorak’s V image: 42′′.51 (± 0′′.05), 39′′.3 (± 0′′.1). Other CCD magnitudes forwarded by the AAVSO for the apparent nova: Nov. 19.27–19.37, 11.6 (A. Oksanen, San Pedro de Atacama, Chile, unfiltered); 19.506, V = 10.17 (W. Dillon, Missouri City, TX, U.S.A.). Visual magnitude estimate by A. Amorim, Florianopolis, Brazil: Nov. 20.036, 9.9.

BRIGHT NOVA IN M31 S. Nakano, Sumoto, Japan, reports the discovery by Koichi Nishiyama

(Kurume, Fukuoka-ken, Japan) and Fujio Kabashima (Miyaki-cho, Sagaken, Japan) of an apparent bright nova in M31 on seven 20-s unfiltered CCD frames taken around Nov. 17.57 UT using a 40-cm reflector. The new star, designated M31N 2007-11d at the CBAT website that catalogues M31

apparent novae (http://www.cfa.harvard.edu/iau/CBAT M31.html) and

located at α = 0^{h}44^{m}54^{s}.60, δ = +41^{o}37′′40′′.0 (equinox 2000.0), was first detected at mag 17.7 on Nov. 16.51 and reached mag 14.9 on Nov. 20.385.

POSSIBLE NOVA IN SERPENS Nakano also reports that Nishiyama and Kabashima have found a vari-

able star located at α = 18^{h}09^{m}24^{s}.25, δ = −7^{o}22′′14′′.2 (equinox 2000.0), with the following unfiltered CCD magnitudes (200-mm camera lens, except 40-cm reflector on Nov. 20): Oct. 2.485 UT, 12.8; 12.501, 12.3; Nov. 19.358, 10.5; 19.387, 10.7; 20.414, 11.2; 20.430, 11.1; 20.431, 11.0. A USNO-B1.0 catalogue star (mag 18.5) has position end figures 24^{s}.30, 13′′.5.

2007 November 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green